1/1

Михаил Юрьевич Лермонтов,  Михаил Юрьевич Лермонтов,
Das Segel  Sail
translated by Heinrich Greif  translated by A. Z. Foreman

Wo Meer und Himmel sich vereinen, Erglänzt ein Segel, weiß und weit - Was trieb es aus dem Land der Seinen? Was sucht es in der Einsamkeit?

A sail drifts white and on its own Amid the light blue ocean haze. What does it seek in distant country? What made it leave its native bays?

Es pfeift der Wind. Die Wellen drohen. Es knarrt der Mast. Das Segel schwebt Nicht vor dem Glück ist es geflohen. Es ist nicht Glück, wonach es strebt.

Great billows play. High winds are whistling Down at the bending, creaking mast Oh! This one seeks no happy ending And does not flee a happy past.

Strahlt auch in Gold der Himmelsbogen, Und glänzt auch noch so blau das Meer - Das Segel lechzt nach Sturm und Wogen, Als ob in Stürmen Ruhe wär.

Beneath, a brighter stream than azure.

Above, the golden sunray flows

Yet this one, restive...quests for tempests As if in tempests were repose.

Translation of the Russian poem “Парус” by Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) into German by Heinrich Greif (∗1907-03-11 – †1946-07-16).

Arg-545-1944 (2015-01-24 19:29:11)

Pri Heinrich Greif vidu la retejon http: // de. wikipedia. org/ wiki/ Heinrich_Greif . La traduko troviĝas en http: // langadvice. blogspot. ru/ 2012/ 10/ blog-post. html .

Translation of the Russian poem “Парус” by Михаил Юрьевич Лермонтов (∗1814-10-16 – †1841-07-27) into English by A. Z. Foreman.

Arg-545-2269 (2015-01-28 19:03:20)

I, Manfred Retzlaff, found this poem-translation in http: // poemsintranslation. blogspot. e/ 2009/ 10/ lermontov-sail-from-russian. html .

http://www.poezio.net/version?poem-id=545&version-id=1944,2269